i didnt know where to ask this question it is for college algebra: solve the system {3x -y+2z=15
x+2y+3z=5
-2x+y-3z=-16

Answer by stanbon(48569)

3x -y+2z=15
x+2y+3z=5
-2x+y-3z=-16
----------
Using matrices and a TI=83 calculator I get:
x=2, y=-3, z= 3
Cheers,
stan H.